OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 7x - 6y = -88.5 -28x + 24y = -354

OpenStudy (anonymous):

No solution First, change the first equation 4(7x - 6y)= 4(-88.5) 28x - 24y = -354 Then, you add or subtract the two equations to eliminate one variable and solve for the other. However... 28x - 24y = -354 -28x + 24y = -354 ------------------ If you add the equations, both variables are eliminated, and if you subtract them, neither equation is eliminated. So, NO SOLUTION

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no prob, post if you have any other qs. & click best response on my post if you think I deserve it!